EDITORIAL

Terry Goodman

Dr. David Ewing, one of my colleagues here at CMSU, is stepping down as our Associate Editor for Mathematics Education for MJMS. David has had primary responsibility for reviewing mathematics education manuscripts submitted to the journal. We appreciate his years of service to MJMS and know that he will continue to make significant contributions in mathematics education. Until we have selected a new Associate Editor to replace Dr. Ewing, I will be reviewing mathematics education manuscripts that we receive.

The No Child Left Behind legislation continues to create much discussion, evaluation, and debate. The results of two “report cards” relative to NCLB can be found at the following addresses:

http://www.nytimes.com/2005/02/24/national/24child.html http://www.fairtest.org/nattest/year%20three%20Report%20Card.htm

President Bush has also proposed a $1.5 billion High School Initiative, which includes a focus on mathematics and science eduction. Specifics in the initiative include the following:

1. Provide a total of $260 million for the Math and Science Partnership program, with $120 million of this funding dedicated to high school mathematics programs.

2. Develop an Adjunct Teacher Corps that would include partnerships between school districts and public/private institutions that would create opportunities for professionals to teach middle and high school courses in core academic courses.

3. Make permanent teacher loan forgiveness in the amount of $17,500 for mathematics, science, and special education teachers who teach at low-income schools.

4. Expand assessments required under NCLB to measure progress with tests in reading and mathematics for grades 9–11.